A cylindrical object weighing 25.0 has the following dimension heights =10.5cm and diameter =2.8 cm calculate the density of this object .
Answer #1

Answer:-

Radius (r) = Diameter / 2 = (2.8 / 2) cm = 1.4 cm

Height (h) = 10.5 cm

Volume of a cylinder = πr​​​​​​​2h = 3.14 × (1.4)2 × 10.5 = 64.6212 cm3

Mass = 25 g **

Density = Mass / Volume = (25 / 64.6212) = 0.387 g/cm3

---------------------------------------------------